# An ordering for the strength of functional dependence

### v1

@@ 5.2 Comparison of $\preccurlyeq_{ccx}$ with the dilation order and dispersive order
$P(Y\geq q_{Y}(W)\mid\mathbb{X}={\bf q}_{\mathbb{X}}(\mathbb{u}))$ (30)
for Lebesgue-almost all $\mathbb{u}\in[0,1]^{p}.$ Here, $W$ is uniform on $(0,1)$ and independent from $(Y,\mathbb{X})\,.$ Similar to the maximum property of the conditional convex order, maximal variability in (30) (in convex order) corresponds to perfect dependence of $Y$ on $\mathbb{X}$; see [22, Proposition 3(ii)]. However, a drawback of this approach is that the random variable in (30) is not constant for independent $\mathbb{X}$ and $Y$. Hence, independence of $\mathbb{X}$ and $Y$ cannot be identified with minimal elements in convex order. Moreover, a comparison of the random variables in (30) with respect to the convex order requires equal means, which is typically not the case for conditional distributions. To overcome the latter shortcoming, the idea in [22] is to center the conditional distributions by subtracting conditional means. This leads to the concept of dilation order defined by
@@
A stronger order, that implies the dilation order, is the dispersive order, also discussed in [22]. The following result can be extended to multivariate vectors $X$ and $X^{\prime}$ with general distribution functions.
Proposition 5.2 (Dilation ordering critertion [22, Theorem 2])
@@
To verify the dilation order in (32), one needs to substract the conditional expectations $\mathbb{E}[F_{Y}(Y)|X=q_{X}(u)]$ and $\mathbb{E}[F_{Y^{\prime}}(Y^{\prime})|X^{\prime}=q_{X^{\prime}}(u)],$ respectively. However, as we discuss in the sequel, the centered conditional distributions are even in the simplest models not comparable in convex order, so that the dilation order in (32) cannot be verified; see Figure 4. To this end, we make use of the following necessary condition for convex order; see [57, Equation (3.A.12)]. We denote by $\mathop{\mathrm{supp}}(S)$ the support of a random variable $S$. Similarly, $\mathop{\mathrm{supp}}(S\mid T=t)$ is the support of the conditional distribution of $S$ given $T=t$.
@@
Consider the additive error models $Y=X+\sigma\varepsilon$ and $Y^{\prime}=X+\sigma^{\prime}\varepsilon$ for $0<\sigma<\sigma^{\prime}$ with $\varepsilon,X$ standard normal and independent. To verify (32), we obtain for $W\sim U(0,1)$ that
$F_{F_{Y}(Y)|X=q_{X}(u)}^{-1}(W)$ $=\Phi\left(\frac{(1+\sigma)\Phi^{-1}(W)-\Phi^{-1}(u)}{\sigma}\right)\quad\text{and}$ (34)
$F_{F_{Y^{\prime}}(Y^{\prime})|X=q_{X}(u)}^{-1}(W)$ $=\Phi\left(\frac{(1+\sigma^{\prime})\Phi^{-1}(W)-\Phi^{-1}(u)}{\sigma^{\prime}}\right)$ (35)
for $\lambda$-almost all $u\in(0,1)\,.$ It is not difficult to see that $\mathbb{E}[F_{Y}(Y)\mid X=q_{X}(u)]\neq\mathbb{E}[F_{Y^{\prime}}(Y^{\prime})\mid X=q_{X}(u)]$ for $\lambda$-almost all $u\in(0,1)\,.$ Since $\mathop{\mathrm{supp}}(F_{Y}(Y)\mid X=q_{X}(u))=\mathop{\mathrm{supp}}(F_{Y^{\prime}}(Y^{\prime})\mid X=q_{X}(u))=[0,1]$ for all $u\in(0,1)\,$, we obtain from Proposition 5.4 for $\lambda$-almost all $u\in(0,1)$ that $(F_{Y}(Y)\mid X=q_{X}(u))$ and $(F_{Y^{\prime}}(Y^{\prime})\mid X^{\prime}=q_{X^{\prime}}(u))$ are not comparable in the dilation order (and thus neither in the dispersive order), see Figure 4. Hence, Proposition 5.2 cannot be applied to prove $\xi(Y,X)\geq\xi(Y^{\prime},X)\,.$ Nevertheless, the inequality for $\xi$ holds as a consequence of Theorem 1.8, where we have shown that $(Y,X)\succcurlyeq_{ccx}(Y^{\prime},X)$.
Figure 4: Left plot: Densities and means of the conditional distributions $F_{Y}(Y)|X=q_{X}(u)$ (solid) and $F_{Y^{\prime}}(Y^{\prime})|X^{\prime}=q_{X^{\prime}}(u)$ (dashed) in Example 5.5 for $\sigma=1$ and $\sigma^{\prime}=2$ and for u=0.1.u=0.1\,. The densities are determined by a kernel estimation based on a sample of size $10^{7}\,.$ Right plot: Densities from the left plot shifted by the mean. Since the supports are non-nested intervals satisfying (33), they are not comparable in convex order.
@@ 5.3 Variability of regression functions and magnitude of error curves
Therefore, let $(Y,X)$ be a bivariate random vector. Then there exists a representation $Y=f(X,U)$ a.s. for a measurable function $f$ and a random variable $U$ that is uniform on $(0,1)$ and independent of $X$. The variable $U$ can be interpreted as noise. Informally, a weak (strong) influence of $U$ on $Y$ indicates strong (weak) dependence of $Y$ on $X$.
The idea in [58] is to define global dependence orderings for the regression function $m(x)=\mathbb{E}[Y|X=x]$ and for the error function $e(x)=\mathrm{Var}(Y|X=x)$. For ordering the magnitude of $e(X)$, the authors propose the (standard) stochastic order. For $m(X)$, they consider variability orderings such as the convex order, the dilation order, and the dispersive order. This approach comes closest to what we have been able to find in the literature on our approach, which compares the variability of the regression functions $P(Y\geq q_{Y}(v)\mid X)=\mathbb{E}[\mathds{1}_{Y\geq q_{Y}(v)}\mid X]$ for all $v\in(0,1)$. This is crucial for a dependence order underlying $\xi$.
@@
Appendix A Auxiliary results and preparation for proofs in Appendix B
@@ A.1 Generalized inverses
    $\iota_{G}(t)=t=\iota_{G}^{-}(t)$ if and only if $G$ is continuous at $G^{-1}(t)$;
@@
    $\iota_{G}^{-}$ is left-continuous while, in general, $\iota_{G}$ is neither left- nor right-continuous;
@@
    $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$;
@@
    $\overline{F}_{Y}(q_{Y}(v))=\overline{F}_{Y^{\prime}}(q_{Y^{\prime}}(v))$ for all $v\in(0,1)$;
@@ A.2 Convex order
For the proof of Theorem 1.4, we use of the following properties of convex order; see [47, Example 1.10.5] and [57, Section 3.A.2].
@@
    Invariance under transformations: $S\leq_{cx}T$ implies $c\,S\leq_{cx}c\,T$ and $S+c\leq_{cx}T+c$ for all constants $c\in\mathbb{R}$.
@@
The conditional convex order is defined by comparing conditional survival probabilities $P(Y\geq q_{Y}(v)\mid\mathbb{X})$ in convex order for all $v\in(0,1)$. Equivalent versions can be achieved for strict survival probabilities and cumulative probabilities as follows. Due to continuity properties of the transformations $\iota_{G}$ and $\iota_{G}^{-1}$ in Proposition A.1(ii) and (iii), exceptional null sets have to be considered.
@@
The following example shows that the exceptional null set in the version of Proposition A.4(iii) is necessary.
@@
Consider a random variable $Y$ with distribution function given by the left-hand graph in Figure 5. Let $Y^{\prime}=Y+2$ if $Y\leq 0.5\,$, and $Y^{\prime}=Y+1.5$ if Y>0.5.Y>0.5\,. Then, for $v=F_{Y}(0.5)=0.4\,$, we have $P(Y\leq q_{Y}(v))=0.4\neq 0.8=P(Y^{\prime}\leq q_{Y^{\prime}}(v))$. Hence, $P(Y\leq q_{Y}(v)\mid\mathbb{X})$ and $P(Y^{\prime}\leq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})$ are not comparable in convex order, independent of the choice of $\mathbb{X}$. However, since $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$ and since $F_{Y}^{-}\circ q_{Y}$ as well as $F_{Y^{\prime}}^{-}\circ q_{Y^{\prime}}$ are left-continuous, it follows that $P(Y\geq q_{Y}(v))=P(Y^{\prime}\geq q_{Y^{\prime}}(v))$ for all $v\in(0,1)$. Consequently, $(Y,\mathbb{X})$ and $(Y^{\prime},\mathbb{X})$ are comparable in conditionally convex order; more precisely, it is $(Y,\mathbb{X})=_{ccx}(F_{Y}(Y),\mathbb{X})=_{ccx}(F_{Y^{\prime}}(Y^{\prime}),\mathbb{X})=_{ccx}(Y^{\prime},\mathbb{X})$ for any choice of $\mathbb{X}$ using distributional invariance of $\preccurlyeq_{ccx}$ and $F_{Y^{\prime}}(Y^{\prime})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}F_{Y}(Y)$.
@@ A.3 Distributional transform and quantile transform
Let $\mathbb{V}=(V_{1},\ldots,V_{p})$ be a random vector (on $(\Omega,\mathcal{A},P)$ which is assumed to be non-atomic) that is independent of $\mathbb{X}=(X_{1},\ldots,X_{p})$ and uniformly on $(0,1)^{p}$ distributed. Then the multivariate distributional transform $\tau_{\mathbb{X}}(\mathbb{X},\mathbb{V})$ of $\mathbb{X}$ (also known as generalized Rosenblatt transform) is defined by
@@
for $\mathbb{x}=(x_{1},\ldots,x_{p})\in\mathbb{R}^{p}$ and all $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{p})\in[0,1]^{p}$, where
@@
see [52, Section 1.3]. For $p=1$ and for a random variable $X$ with continuous distribution function $F_{X}\,$, the distributional transform $\tau_{X}(X,V)$ simplifies to $F_{X}(X)\,$, which is uniform on $(0,1)$.
@@
$\begin{split}\zeta_{1}&:=F_{X_{1}}^{-1}(U_{1})\,,\\ \zeta_{i}&:=F^{-1}_{X_{i}|X_{i-1}=\xi_{i-1},\ldots,X_{1}=\xi_{1}}(U_{i})\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{for all }i\in\{2,\ldots,p\}\end{split}$ (36)
@@
${\bf q}_{\mathbb{X}}(\mathbb{U})$ $\text{ is a random vector with distribution function }F_{\mathbb{X}},$ (38)
@@
$\mathbb{X}={\bf q}_{\mathbb{X}}\big(\tau_{\mathbb{X}}(\mathbb{X},\mathbb{V})\big)\qquad P\text{-almost surely.}$ (39)
Appendix B Proofs
@@
The random variable $P(Y\geq q_{Y}(v)\mid\mathbb{X})$ is a maximal element in convex order within the class of $[0,1]$-valued random variables having mean $P(Y\geq q_{Y}(v))$; see Lemma A.3. Hence, for $(Y^{\prime}\mathbb{X}^{\prime})\in\mathcal{R}_{A}$, using the marginal constraint, it follows that $P(Y^{\prime}\geq q_{Y^{\prime}}(v))=P(Y\geq q_{Y}(v))$ and thus $P(Y^{\prime}\geq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})\leq_{cx}P(Y\geq q_{Y}(v)\mid\mathbb{X})$ for all $v\in(0,1)$. Therefore, $(Y^{\prime},\mathbb{X}^{\prime})\preccurlyeq_{ccx}(Y,\mathbb{X})$ for all random vectors $(Y^{\prime},\mathbb{X}^{\prime})\in\mathcal{R}_{A}$. For the reverse direction, assume that $(Y^{\prime},\mathbb{X}^{\prime})\preccurlyeq_{ccx}(Y,\mathbb{X})$ for all $(Y^{\prime},\mathbb{X}^{\prime})\in\mathcal{R}_{A}$. This implies, in particular, $(Y,Y)\preccurlyeq_{ccx}(Y,\mathbb{X})$ and thus for all $v\in(0,1)$ that
@@
To verify Axiom 7, observe that $\overline{\mathsf{Ran}(F_{g(Y)})}=\overline{\mathsf{Ran}(F_{Y})}$. It follows that $P(g(Y)\geq q_{g(Y)}(v))=P(Y\geq q_{Y}(v))$ for all $v\in(0,1)\,;$ see Lemma A.2. Hence, the invariance property for $Y$ follows from the definition of conditional probability and from the definition of conditional convex order. To show the invariance property for $\mathbb{X}$, let $\mathbb{h}$ be bijective. This implies that the $\sigma$-algebras generated by $\mathbb{X}$ and $\mathbb{h}(\mathbb{X})$ coincide. Hence, the statement follows from the definition of the conditional convex order.
@@
The proof of Theorem 1.6 is based on Theorem 2.2 and Proposition 2.7 which we prove first.
@@
For the proof of Proposition 2.7, we make use of the following result on rearrangements of conditional distribution functions.
@@
To show statement (i), we obtain for all $y\in\mathbb{R}$ that
$P\big(q_{Y|\mathbb{X}}^{\uparrow U}(V)\leq y\big)$ $=\int_{[0,1]}P(F_{u}^{-1}(V)\leq y\mid U=u)\mathsf{\,d}\lambda(u)=\int_{[0,1]}P(F_{u}^{-1}(V)\leq y)\mathsf{\,d}\lambda(u)$
@@
where the use for the first equality the definition of the rearranged quantile transform in (10). For the second equality, we use independence of $U$ and $V$. The third equality follows from Proposition A.1(i), using the fact that $F_{u}$ is a distribution function, and from the fact that $V$ is uniform on $(0,1)$. The fourth equality holds by definition of $F_{u}$ as a rearrangement of $\mathbb{u}\mapsto P(Y\leq y\mid\mathbb{X}=q_{\mathbb{X}}(\mathbb{u}))$. The fifth equality follows with (38), and the last equality is due to disintegration.
To show statement (ii), recall that the family $\{q_{Y|\mathbb{X}}^{\uparrow u}(V)\}_{[0,1]}$ is SI in $u$, that is, $\mathbb{E}f(q_{Y|\mathbb{X}}^{\uparrow u}(V))=\mathbb{E}[f(q_{Y|\mathbb{X}}^{\uparrow u}(V))\mid U=u]$ is increasing in $u$ for all increasing functions $f$ such that the expectations exist, where we use for the equality that $U$ and $V$ are independent. However, the latter means that $(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ is a bivariate SI random vector.
To show statement (iii), let $v\in(0,1)$. Then we obtain for all $u\in[0,1]$ that
$\int_{0}^{u}P\big(q_{Y|\mathbb{X}}^{\uparrow U}(V)\leq q_{q_{Y|\mathbb{X}}^{\uparrow U}}(v)\mid U=s\big)\mathsf{\,d}s$ $=\int_{0}^{u}P\big(q_{Y|\mathbb{X}}^{\uparrow s}(V)\leq q_{Y}(v)\big)\mathsf{\,d}s$
@@
where we use for the first equality $q_{Y|\mathbb{X}}^{\uparrow U}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y$ and independence of $U$ and $V$. The second equality follows from (10) and Proposition A.1(i). The last equality is due to (9). Since the integrand of the first integral is decreasing in $s$, it coincides $\lambda$-almost surely with the decreasing rearrangement $(\eta_{Y|\mathbb{X}}^{v})^{*}$. Hence, the characterization of $\preccurlyeq_{ccx}$ through the Schur order in Theorem 2.2 implies $(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=_{ccx}(Y,\mathbb{X})$. $\blacksquare$

### v2

@@ 5.2 Comparison of $\preccurlyeq_{ccx}$ with the dilation order and dispersive order
$P(Y\geq q_{Y}(V)\mid\mathbb{X}={\bf q}_{\mathbb{X}}(\mathbb{u}))$ (40)
for Lebesgue-almost all $\mathbb{u}\in[0,1]^{p}.$ Here, $V\sim U(0,1)$ is independent from $\mathbb{X}$. Similar as for $\preccurlyeq_{ccx}$, maximal variability in (40) (in convex order) corresponds to perfect dependence of $Y$ on $\mathbb{X}$; see [22, Proposition 3 (ii)]. However, a drawback of this approach is that the random variable in (40) is not constant for independent $\mathbb{X}$ and $Y$. Hence, independence of $\mathbb{X}$ and $Y$ cannot be identified with minimal elements in convex order. Moreover, a comparison of the random variables in (40) with respect to the convex order requires equal means, which is typically not the case for conditional distributions. To overcome the latter shortcoming, the idea in [22] is to center the conditional distributions by subtracting conditional means. This leads to the concept of dilation order defined by
@@
A stronger order, that implies the dilation order, is the dispersive order, also discussed in [22]. The following result can be extended to arbitrary random vectors $\mathbb{X}$ and $\mathbb{X}^{\prime}$.
Proposition 5.2 (Dilation order criterion [22, Theorem 2])
@@
To verify the dilation order in (42), one needs to substract the conditional expectations $\mathbb{E}[F_{Y}(Y)|X=q_{X}(u)]$ and $\mathbb{E}[F_{Y^{\prime}}(Y^{\prime})|X^{\prime}=q_{X^{\prime}}(u)],$ respectively. However, as we discuss in the sequel, the centered conditional distributions are even in the simplest models not comparable in convex order, so that the dilation order in (42) cannot be verified; see Figure 4. To this end, we use the following necessary condition for convex order; see [58, Equation (3.A.12)]. We denote by $\mathop{\mathrm{supp}}(S)$ the support of a random variable $S$. Similarly, $\mathop{\mathrm{supp}}(S\mid T=t)$ is the support of the conditional distribution of $S$ given $T=t$.
@@
Consider the additive error models $Y=X+\sigma\varepsilon$ and $Y^{\prime}=X+\sigma^{\prime}\varepsilon$ for $0<\sigma<\sigma^{\prime}$ where $\varepsilon,X$ are standard normal and independent. To verify (42), we obtain for $V\sim U(0,1)$ independent of $X$ that
$F_{F_{Y}(Y)|X=q_{X}(u)}^{-1}(V)$ $=\Phi\left(\frac{(1+\sigma)\Phi^{-1}(V)-\Phi^{-1}(u)}{\sigma}\right)\quad\text{and}$ (44)
$F_{F_{Y^{\prime}}(Y^{\prime})|V=q_{X}(u)}^{-1}(V)$ $=\Phi\left(\frac{(1+\sigma^{\prime})\Phi^{-1}(V)-\Phi^{-1}(u)}{\sigma^{\prime}}\right)$ (45)
for $\lambda$-almost all $u\in(0,1)\,.$ It is not difficult to see that $\mathbb{E}[F_{Y}(Y)\mid X=q_{X}(u)]\neq\mathbb{E}[F_{Y^{\prime}}(Y^{\prime})\mid X=q_{X}(u)]$ for $\lambda$-almost all $u\in(0,1)\,.$ Since $\mathop{\mathrm{supp}}(F_{Y}(Y)\mid X=q_{X}(u))=\mathop{\mathrm{supp}}(F_{Y^{\prime}}(Y^{\prime})\mid X=q_{X}(u))=[0,1]$ for all $u\in(0,1)\,$, we obtain from Proposition 5.4 for $\lambda$-almost all $u\in(0,1)$ that $(F_{Y}(Y)\mid X=q_{X}(u))$ and $(F_{Y^{\prime}}(Y^{\prime})\mid X^{\prime}=q_{X^{\prime}}(u))$ are not comparable in the dilation order (and thus neither in the dispersive order), see Figure 4. Hence, Proposition 5.2 cannot be applied to prove $\xi(Y,X)\geq\xi(Y^{\prime},X)\,.$ Nonetheless, this inequality for $\xi$ follows from Theorem 1.8 where we have shown that $(Y,X)\succcurlyeq_{ccx}(Y^{\prime},X)$.
Figure 4: Left plot: Densities and means of the conditional distributions $F_{Y}(Y)|X=q_{X}(0.1)$ for $\sigma=1$ (dashed) and $\sigma=2$ (solid). Right plot: Densities from the left plot shifted by the mean. Since the supports are non-nested intervals satisfying (43), they are not comparable in convex order.
@@ 5.3 Variability of regression functions and magnitude of error curves
Therefore, let $(Y,X)$ be a bivariate random vector. Then there exists a representation $Y=f(X,U)$ a.s. for a measurable function $f$ and a random variable $U$ that is uniform on $(0,1)$ and independent of $X$. The random variable $U$ can be interpreted as noise. Informally, a weak (strong) influence of $U$ on $Y$ indicates strong (weak) dependence of $Y$ on $X$.
The idea in [59] is to define global dependence orderings for the regression function $m(x)=\mathbb{E}[Y|X=x]$ and for the error function $e(x)=\mathrm{Var}(Y|X=x)$. For ordering the magnitude of $e(X)$, the authors propose the (standard) stochastic order. For $m(X)$, they consider variability orderings such as the convex order, the dilation order, and the dispersive order. This approach comes closest to what we have been able to find in the literature on our approach, which compares the variability of the regression functions $P(Y\geq q_{Y}(v)\mid X)=\mathbb{E}[\mathds{1}_{\{\{Y\geq q_{Y}(v)\}\}}\mid X]$ for all $v\in(0,1)$. This is crucial for a dependence order underlying $\xi$.
@@
A Auxiliary results and preparation for proofs in Appendix B
@@ A.1 Generalized inverses
    $G^{-1}\circ G\circ G^{-1}(t)=G^{-1}(t)$ for all $t\in(0,1)$,
@@
    $\iota_{G}(t)=t=\iota_{G}^{-}(t)$ if and only if $G$ is continuous at $G^{-1}(t)$,
@@
    $\iota_{G}^{-}$ is left-continuous while, in general, $\iota_{G}$ is neither left- nor right-continuous,
  5. (v)
@@
    $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$,
@@
    $\overline{F}_{Y}(q_{Y}(v))=\overline{F}_{Y^{\prime}}(q_{Y^{\prime}}(v))$ for all $v\in(0,1)$,
@@ A.2 Convex order
For the proofs of several results, we use the following properties of convex order; see [45, Example 1.10.5] and [58, Section 3.A.2].
@@
    Invariance under linear transformations: $S\leq_{cx}T$ implies $c\,S+d\leq_{cx}c\,T+d$ for all constants $c,d\in\mathbb{R}$.
@@
The conditional convex order is defined by comparing conditional survival probabilities $P(Y\geq q_{Y}(v)\mid\mathbb{X})$ in convex order for all $v\in(0,1)$. Equivalent versions can be achieved for strict survival probabilities and cumulative probabilities as follows. Due to continuity properties of the transformations $\iota_{G}$ and $\iota_{G}^{-}$ in Proposition A.1iii and iv, exceptional null sets have to be considered.
@@
The following example shows that the exceptional null set in the version of Proposition A.4iii is necessary.
@@
Consider a random variable $Y$ with distribution function given by the left-hand graph in Figure 5. The right-hand graph is the distribution function of the random variable $Y^{\prime}=Y+2$ if $Y\leq 0.5\,$, and $Y^{\prime}=Y+1.5$ if Y>0.5.Y>0.5\,. We observe that, for $v=F_{Y}(0.5)=0.4$, it is $P(Y\leq q_{Y}(v))=0.4\neq 0.8=P(Y^{\prime}\leq q_{Y^{\prime}}(v))$. Hence, $P(Y\leq q_{Y}(v)\mid\mathbb{X})$ and $P(Y^{\prime}\leq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})$ are not comparable in convex order, independent of the choice of $\mathbb{X}$. However, since $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}$ and since $F_{Y}^{-}\circ q_{Y}$ and $F_{Y^{\prime}}^{-}\circ q_{Y^{\prime}}$ are left-continuous, it follows that $P(Y\geq q_{Y}(v))=P(Y^{\prime}\geq q_{Y^{\prime}}(v))$ for all $v\in(0,1)$. Consequently, $(Y,\mathbb{X})$ and $(Y^{\prime},\mathbb{X})$ are comparable in conditional convex order; more precisely, it is $(Y,\mathbb{X})=_{ccx}(F_{Y}(Y),\mathbb{X})=_{ccx}(F_{Y^{\prime}}(Y^{\prime}),\mathbb{X})=_{ccx}(Y^{\prime},\mathbb{X})$ for any choice of $\mathbb{X}$ using distributional invariance of $\preccurlyeq_{ccx}$ and $F_{Y^{\prime}}(Y^{\prime})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}F_{Y}(Y)$.
@@ A.3 Distributional transform and quantile transform
Let $\mathbb{V}=(V_{1},\ldots,V_{p})$ be a random vector that is independent of $\mathbb{X}=(X_{1},\ldots,X_{p})$ and uniform on $(0,1)^{p}$. Then the multivariate distributional transform $\tau_{\mathbb{X}}(\mathbb{X},\mathbb{V})$ of $\mathbb{X}$ (also known as generalized Rosenblatt transform) is defined by
@@
for $\mathbb{x}=(x_{1},\ldots,x_{p})\in\mathbb{R}^{p}$ and $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{p})\in[0,1]^{p}$, where
@@
see [51, Section 1.3]. For $p=1$ and for a random variable $X$ with continuous distribution function $F_{X}\,$, the distributional transform $\tau_{X}(X,V)$ simplifies to $F_{X}(X)\,$, which is uniform on $(0,1)$.
@@
$\begin{split}\zeta_{1}&:=F_{X_{1}}^{-1}(U_{1})\,,\\ \zeta_{i}&:=F^{-1}_{X_{i}|X_{i-1}=\xi_{i-1},\ldots,X_{1}=\xi_{1}}(U_{i})\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{for all }i\in\{2,\ldots,p\}\end{split}$ (A.1)
@@
${\bf q}_{\mathbb{X}}(\mathbb{U})$ $\text{ is a random vector with distribution function }F_{\mathbb{X}},$ (A.3)
@@
$\mathbb{X}={\bf q}_{\mathbb{X}}\big(\tau_{\mathbb{X}}(\mathbb{X},\mathbb{V})\big)\qquad P\text{-almost surely.}$ (A.4)
B Proofs
@@
The random variable $P(Y\geq q_{Y}(v)\mid\mathbb{X})$ is a maximal element in convex order within the class of $[0,1]$-valued random variables having mean $P(Y\geq q_{Y}(v))$; see Lemma A.3. Hence, for $(Y^{\prime},\mathbb{X}^{\prime})\in\mathcal{R}_{A}$, using the marginal constraint, it follows that $P(Y^{\prime}\geq q_{Y^{\prime}}(v))=P(Y\geq q_{Y}(v))$ and thus $P(Y^{\prime}\geq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})\leq_{cx}P(Y\geq q_{Y}(v)\mid\mathbb{X})$ for all $v\in(0,1)$. Therefore, $(Y^{\prime},\mathbb{X}^{\prime})\preccurlyeq_{ccx}(Y,\mathbb{X})$ for all random vectors $(Y^{\prime},\mathbb{X}^{\prime})\in\mathcal{R}_{A}$. For the reverse direction, assume that $(Y^{\prime},\mathbb{X}^{\prime})\preccurlyeq_{ccx}(Y,\mathbb{X})$ for all $(Y^{\prime},\mathbb{X}^{\prime})\in\mathcal{R}_{A}$. This implies, in particular, $(Y,Y)\preccurlyeq_{ccx}(Y,\mathbb{X})$ and thus for all $v\in(0,1)$ that
@@
To verify Axiom 7, observe that $\overline{\mathsf{Ran}(F_{g(Y)})}=\overline{\mathsf{Ran}(F_{Y})}$. It follows that $P(g(Y)\geq q_{g(Y)}(v))=P(Y\geq q_{Y}(v))$ for all $v\in(0,1)\,;$ see Lemma A.2. Hence, the invariance property for $Y$ follows from the definition of conditional convex order. To show the invariance property for $\mathbb{X}$, let $\mathbb{h}$ be bijective. This implies that the $\sigma$-algebras generated by $\mathbb{X}$ and $\mathbb{h}(\mathbb{X})$ coincide. Hence, the statement follows from the definition of the conditional convex order.
@@
The proof of Theorem 1.6 is based on several auxiliary results as well as on Theorem 2.2 and Theorem 2.7 which we prove first.
@@
For the proof of Theorem 2.7, we make use of the following result on rearrangements of conditional distribution functions.
@@
To show statement i, we obtain for all $y\in\mathbb{R}$ that
$P\big(q_{Y;\mathbb{X}}^{\uparrow U}(V)\leq y\big)$ $=\int_{[0,1]}P(F_{u}^{-1}(V)\leq y\mid U=u)\mathsf{\,d}\lambda(u)=\int_{[0,1]}P(F_{u}^{-1}(V)\leq y)\mathsf{\,d}\lambda(u)$
@@
where the use for the first equality the definition of the rearranged quantile transform in (13). For the second equality, we use independence of $U$ and $V$. The third equality follows from Proposition A.1i, using the fact that $F_{u}$ is a distribution function, and from the fact that $V$ is uniform on $(0,1)$. The fourth equality holds by definition of $F_{u}$ as a rearrangement of $\mathbb{u}\mapsto P(Y\leq y\mid\mathbb{X}=q_{\mathbb{X}}(\mathbb{u}))$. The fifth equality follows with (A.3), and the last equality is due to disintegration.
To show statement ii, recall that the family $\{q_{Y;\mathbb{X}}^{\uparrow u}(V)\}_{[0,1]}$ is SI in $u$, that is, $\mathbb{E}f(q_{Y;\mathbb{X}}^{\uparrow u}(V))=\mathbb{E}[f(q_{Y;\mathbb{X}}^{\uparrow u}(V))\mid U=u]$ is increasing in $u$ for all increasing functions $f$ such that the expectations exist, where we use for the equality that $U$ and $V$ are independent. However, the latter means that $(q_{Y;\mathbb{X}}^{\uparrow U}(V),U)$ is a bivariate SI random vector.
To show statement iii, let $v\in(0,1)$. Then we obtain for all $u\in[0,1]$ that
$\int_{0}^{u}P\big(q_{Y;\mathbb{X}}^{\uparrow U}(V)\leq q_{q_{Y;\mathbb{X}}^{\uparrow U}}(v)\mid U=s\big)\mathsf{\,d}s$ $=\int_{0}^{u}P\big(q_{Y;\mathbb{X}}^{\uparrow s}(V)\leq q_{Y}(v)\big)\mathsf{\,d}s$
@@
where we use for the first equality $q_{Y;\mathbb{X}}^{\uparrow U}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y$ and independence of $U$ and $V$. The second equality follows from (13) and Proposition A.1i. The last equality is due to (12). Since the integrand of the first integral is decreasing in $s$, it coincides $\lambda$-almost surely with the decreasing rearrangement $(\eta_{Y|\mathbb{X}}^{v})^{*}$. Hence, the characterization of $\preccurlyeq_{ccx}$ through the Schur order in Theorem 2.2 implies $(q_{Y;\mathbb{X}}^{\uparrow U}(V),U)=_{ccx}(Y,\mathbb{X})$. $\blacksquare$